OpenStudy (anonymous):

Solve the formula for the area of a triangle A = (1/2) bh for height h.

OpenStudy (anonymous):

pic?

OpenStudy (anonymous):

there wasnt one but theres a second question pertaining to the first question

OpenStudy (anonymous):

Using the above formula find h when A = 24 and b = 8.

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

sub in your values given A = 1/2 * b * h 24 = 1/2 * 8 * h 24 = 4h 24/4 = h 6 = h any questions at all ?

OpenStudy (anonymous):

just divide 24 by 4

OpenStudy (anonymous):

yea id get bad points for not showing work

OpenStudy (texaschic101):

show the entire process so the asker will understand...do not tell him to just divide 24 by 4

OpenStudy (anonymous):

and thank you oh and did you answer number one or number two

OpenStudy (texaschic101):

I found h when A = 24 and b = 8.....h = 6

OpenStudy (anonymous):

ok that was number two can you help with number one plsssss

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

We have to do inverse operations to get something by itself. To get h by itself you have to ask what is happening to h. h is being multiplied by b and multiplied by 1/2 The inverse divide by b and divide by 1/2 Then h is by itself. A=1/2bh divide by b A/b=1/2h divide by 1/2 (which is the same as multiplying 2/1 remember the rules from fractions) 2A/b=h

OpenStudy (anonymous):

I hope this help! :D

OpenStudy (texaschic101):

A = 1/2bh....looking for h multiply both sides by 2 2A = bh -- now divide both sides by b (2A)/b = h I agree with EmoBae ^^